# GMODiff: One-Step Gain Map Refinement with Diffusion Priors for HDR Reconstruction

### v1

@@ 4.3 Ablation Studies
Effects of HDR Reconstruction Loss. During the training of GMODiff, we replace the logarithmic compression in Eq. (1), which is non-invertible at zero, with $\mu$-law companding to enable direct supervision using ground-truth HDR images. This modification allows us to reconstruct the HDR image directly from the estimated gain map and apply pixel-wise supervision in the linear HDR domain. As shown in Tab. 2, ablating this reconstruction loss leads to clear performance degradation: minor inaccuracies in gain map estimation are amplified during HDR reconstruction, resulting in reduced fidelity and validating its importance for stable and accurate HDR reconstruction.
Effects of Degradation-Aware Priors. The degradation-aware priors, derived from intermediate features of the first-stage regression branch that predicts a degradation mask, implicitly encode spatially varying artifact patterns. Although the mask itself is not used as direct conditioning, its underlying feature representation provides useful cues about input-specific degradation. As shown in Tab. 2, removing these priors leads to a consistent drop across all metrics-both perceptual (e.g., LPIPS, MUSIQ) and distortion-based (e.g., PU-PSNR). This demonstrates that even implicit access to degradation-related features benefits HDR reconstruction, enabling the diffusion model to better adapt to input artifacts and produce more accurate and perceptually plausible results.
Effects of Group Training Strategy. To prevent DaReg from overfitting to the training set, which could hinder GMODiff’s ability to effectively learn artifact removal, we propose a Group Training strategy. As illustrated in Fig. 6, without this strategy, GMODiff can still mitigate degradation in the initial gain map estimation and perform basic artifact correction. However, when Group Training is applied, its artifact removal capability is significantly enhanced, yielding results with nearly imperceptible artifacts. Tab. 2 shows improved perceptual scores with Group Training, while distortion metrics slightly decrease, indicating that Group Training achieves a favorable trade-off between fidelity and perceptual quality.
Effects of DA Decoder. As shown in Tab. 2, removing the DA Decoder leads to a significant drop across all evaluation metrics, especially reference-based ones. This degradation stems from the highly compressed latent space of LDMs, which struggles to preserve fine details. To mitigate this limitation, we inject high-fidelity detail features extracted from intermediate layers of the gain map encoder and selectively refine them using the regression prior obtained from the first stage. As illustrated in Fig. 6, the DA Decoder effectively enhances texture fidelity while suppressing artifacts and ghosting present in the initial gain map. This improvement underscores the critical role of the DA Decoder in recovering high-quality details during HDR reconstruction, thereby substantially boosting the perceptual quality of the final output.
Table 2: Ablation study of the each component. The best and second best results are colored with red and blue, respectively.
@@
DaReg 41.67 0.0432 0.5803 59.77
w/o $\mathcal{L}_{r}$ 41.54 0.0415 0.5878 60.74
w/o Group Training 42.19 0.0353 0.5848 60.65
w/o $c_{L}$ 41.76 0.0389 0.5859 60.59
w/o DA Decoder 35.78 0.1076 0.5763 59.89
GMODiff (ours) 41.95 0.0346 0.5884 60.81
Table 3: Inference time comparison on a $1080\times 1920$ image.
Time (s) $\downarrow$ AHDR DiffHDR AFUNet SCTNet UltraFusion Ours
Inference Time 0.24s 13.52s 6.16s 5.66s 93.4s 0.89s
@@ 4.4 Inference Time Analysis
Tab. 3 compares inference latency on $1920\times 1080$ HDR reconstruction. Our method is overall second-fastest, slightly slower than the lightweight CNN-based AHDRNet [yan2019attention], and significantly faster than transformer-heavy models (AFUNet [li2025afunet], SCTNet [tel2023alignment]), which, despite using efficient window-based attention, incur high latency due to deep stacking of Transformer blocks and processing at high spatial resolutions. Among diffusion-based approaches, UltraFusion [chen2025ultrafusion] requires over a minute per image, owing to patch-wise processing with 50 denoising steps, optical flow alignment, and ControlNet overhead. DiffHDR [yan2023toward], despite using only 5 steps, operates in pixel space without latent compression, resulting in high memory bandwidth and slow inference. In contrast, our method leverages a pre-trained LDM that performs computation in a highly compressed latent space with a single denoising step, achieving the fastest speed among diffusion-based methods. Moreover, it is compatible with efficient inference frameworks such as xFormers, which further reduce latency through memory, efficient attention and kernel fusion, making it practical for real-world applications.
Figure 6: Qualitative results of our ablation study.

### v2

@@ 4.3 Ablation Studies
To validate the effectiveness of each core component in the proposed GMODiff framework, we conduct comprehensive ablation experiments on the benchmark test set, with quantitative results reported in Table LABEL:abb and qualitative results presented in Fig. 7. We take the full GMODiff model as the baseline to ablate key modules one by one, and also evaluate the standalone performance of DaReg for comparison. DaReg-Seg refers to the joint training of the reconstruction branch and segmentation head, which leads to a noticeable performance drop. In the second stage of the framework, removing the DA Decoder causes a significant performance decline, as the latent space of LDM fails to preserve the rich detail information contained in the initial GM. Thanks to the rational design of the overall framework, the model only experiences slight performance degradation when other modules are ablated. Compared with the original regression-based DaReg model, GMODiff achieves a favorable trade-off between perceptual quality and distortion metrics with acceptable computational overhead.
Figure 7: Qualitative results of our ablation study.
Table 3: Inference time comparison on a $1080\times 1920$ image.
Time (s) $\downarrow$ AHDR [yan2019attention] DiffHDR [yan2023toward] AFUNet [kong2024safnet] SCTNet [tel2023alignment] UltraFusion [chen2025ultrafusion] GMDD [guan2025hdr] LE-Diff [wang2025lediff] Ours
Inference Time 0.24s 13.52s 6.16s 5.66s 93.4s 39.2s 132.7s 0.93s
Figure 8: Ablation study of the each component. The best and second best results are colored with red and blue.
@@
DaReg 36.87 0.5819 60.57
DaReg-Seg 35.47 0.5803 59.77
w/o $\mathcal{L}_{r}$ 36.60 0.5875 60.84
w/o Group Training 36.76 0.5858 60.75
w/o DA Mask 36.49 0.5862 60.73
w/o DA Decoder 30.57 0.5754 59.93
GMODiff (ours) 36.68 0.5881 60.96
Figure 9: The user study results on the benchmark test set. Details can be found in the supplementary materials.
@@ 4.4 Inference Time Analysis
Tab. 3 compares inference latency on $1920\times 1080$ HDR reconstruction. Our method is overall second-fastest, slightly slower than the lightweight CNN-based AHDRNet [yan2019attention], and significantly faster than transformer-heavy models (AFUNet [li2025afunet], SCTNet [tel2023alignment]), which, despite using efficient window-based attention, incur high latency due to deep stacking of Transformer blocks and processing at high spatial resolutions. Among diffusion-based approaches, UltraFusion [chen2025ultrafusion] requires over a minute per image, owing to patch-wise processing with 50 denoising steps, optical flow alignment, and ControlNet overhead. DiffHDR [yan2023toward], despite using only 5 steps, operates in pixel space without latent compression, resulting in high memory bandwidth and slow inference. Since LE-Diff [wang2025lediff] contains two LDMs that separately predict results for different exposures, it takes more than 2 minutes to infer one complete HDR image. In contrast, our method leverages a pre-trained LDM that performs computation in a highly compressed latent space with a single denoising step, achieving the fastest speed among diffusion-based methods. Moreover, it is compatible with efficient inference frameworks such as xFormers, which further reduce latency through memory, efficient attention and kernel fusion, making it practical for real-world applications.